# TinySQL: A Progressive Text-to-SQL Dataset for Mechanistic Interpretability Research

### v4

@@ Abstract
Abir Harrasse*1,5  Philip Quirke*1thanks: * Equal contributions. Correspondence to: philip@withmartian.com.  Clement Neo*2 Dhruv Nathawani3Luke Marks1Amir Abdullah4,1 1Martian  2Apart Research  3NVIDIA  4Thoughtworks  5Mohammed VI Polytechnic University
@@ 1 Introduction
Figure 1: (a) TinySQL is broken down into 7 subsets of varying complexities, across both SQL query and user query axes. (b) We train and release a comprehensive set of models on each dataset subset.(c) We apply MI techniques across various configurations to understand model behavior and compare results.
@@
 [^1] Code available at: https://github.com/withmartian/TinySQL,
Datasets, Models and SAEs at: https://huggingface.co/collections/withmartian/tinysql-6760e92748b63fa56a6ffc9f
@@ 4.3 Finding Locally Minimal Text-to-SQL Circuits via EAP and Ablation
For each feature, we ran EAP four times across all combinations of semantic field and table name variations. While we obtain 10 edges per batch, we retain only the edges that appear across batches consistently - requiring 100% consistency for CS1 and CS2, and 80% for CS3 due to its increased complexity. The full results appear in Figures 14, 15, and 16 in the appendix. After running EAP for all features, we obtain a set of edges per feature. We then take the set of attention heads associated with these edges across all features, and retain them during the following ablation study.
@@ 4.4 Sparse Circuit Identification: Method and Results
We begin by collecting attention heads and MLP activations from each of $\texttt{BM}_{2}\_\texttt{CS}_{1}\_\texttt{Syn}$ and $\texttt{BM}_{1}\_\texttt{CS}_{i}\_\texttt{Syn}$ models when applied to its corresponding $\texttt{CS}_{i}$ datasets, for $i\in\{1,3,4,5\}$ To analyze how models trained on more complex tasks generalize to simpler ones, we also create a descending complexity evaluation set by applying BM1_CS3_Syn to CS1 and CS2 data.
[…]
@@ 5.1 Identifiability via SAE-EAP Overlap
In the following, we always use Synonyms circuits so we will omit the notation SYN.
@@ 5.2 Local Minimality of Circuits
We also analyze the local minimality of the circuits—(i.e., the property that no smaller circuit can be obtained by removing or editing components of the current “locally minimal") circuits identified by SAEs, by measuring the percentage of the total model they utilize. This metric reflects how sparse and efficient the discovered subcircuits are in reproducing task behavior. Table 5 summarizes these results across different combinations of models and datasets.
@@
For example, in BM1-CS3 $\rightarrow$ CS1, the found circuit recovered 88.63% accuracy while using only 22.5% of model components. Other cases show similar patterns, with best circuits retaining high performance using only  20–30% of the model.
Finally, on BM2-CS1 $\rightarrow$ CS1, focusing only on attention heads yielded circuits using just 3.17% of the model’s heads while maintaining high accuracy. However, including MLPs raised usage to 60.71%, indicating that while attention sparsity is possible, full circuit function often relies on broader model interaction.
@@ 6 Two-Phase SQL Generation: Intent First, Grounding Later
Figure 3: Emergence of SQL intent (via keywords) precedes table name resolution during generation in BM2CS3. LogitLens reveals elevated logit probabilities for SQL keywords early, followed by table tokens at deeper layers.
@@ G.1.2 EAP results
For each run, we select the top 10 edges per batch and retain only the edges that appeared at least Tof the time across batches. For CS1 and CS2, we set T = 100, while for CS3, we set T = 80.

### v5

@@ Abstract
***Proceedings of the 2025 Conference on Empirical Methods in Natural Language Processing (EMNLP 2025)
@@
Abir Harrasse*1,5   Philip Quirke*1   Clement Neo2 Dhruv Nathawani3Luke Marks1Amir Abdullah4,1 1Martian  2Apart Research  3NVIDIA 4Thoughtworks  5Mohammed VI Polytechnic University *Equal contribution Correspondence to: {philip@withmartian.com, abirharrasse@gmail.com}
@@ 1 Introduction
Figure 1: (a) TinySQL is broken down into 9 subsets of varying complexities, across both SQL query and user query axes. (b) We train and release a comprehensive set of models on each dataset subset.(c) We apply MI techniques across various configurations to understand model behavior and compare results.
@@
 [^2] Project website with code, data, models and viewers at: https://abirharrasse.github.io/tinysql/
@@ 4.3 Finding Locally Minimal Text-to-SQL Circuits via EAP and Ablation
For each feature, we ran EAP four times across all combinations of semantic field and table name variations. While we obtain 10 edges per batch, we retain only the edges that appear across batches consistently, requiring 100% consistency for CS1 and CS2, and 80% for CS3 due to its increased complexity. The full results appear in Figures 14, 15, and 16 in the appendix. After running EAP for all features, we obtain a set of edges per feature. We then take the set of attention heads associated with these edges across all features, and retain them during the following ablation study.
@@ 4.4 Sparse Circuit Identification: Method and Results
We begin by collecting attention heads and MLP activations from each of $\texttt{BM}_{2}\_\texttt{CS}_{1}\_\texttt{Syn}$ and $\texttt{BM}_{1}\_\texttt{CS}_{i}\_\texttt{Syn}$ models when applied to its corresponding $\texttt{CS}_{i}$ datasets, for $i\in\{1,3,4,5\}$ . To analyze how models trained on more complex tasks generalize to simpler ones, we also create a descending complexity evaluation set by applying BM1_CS3_Syn to CS1 and CS2 data.
We train SAEs on the collected activations using the Belrose (2024)’s package, following the top-$k$ training regime Gao et al. (2024) , with $k=16$, an expansion factor of 4, and a learning rate of $8\times 10^{-4}$ for 6 training epochs. Across all models and datasets, we find that the Fraction of Unexplained Variance (FVU) remains below 0.05, indicating that the SAEs faithfully reconstruct the input representations. Training each SAE takes approximately 1 hour for BM1 and 2 hours for BM2, all on a single A100 GPU.
@@ 5.2 Local Minimality of Circuits
We also analyze the local minimality of the circuits (i.e., the property that no smaller circuit can be obtained by removing or editing components of the current “locally minimal") circuits identified by SAEs, by measuring the percentage of the total model they utilize. This metric reflects how sparse and efficient the discovered subcircuits are in reproducing task behavior. Table 5 summarizes these results across different combinations of models and datasets.
@@
For example, in BM1-CS3-Syn $\rightarrow$ CS1, the found circuit recovered 88.63% accuracy while using only 22.5% of model components. Other cases show similar patterns, with best circuits retaining high performance using only  20–30% of the model.
Finally, on BM2-CS1-Syn $\rightarrow$ CS1, focusing only on attention heads yielded circuits using just 3.17% of the model’s heads while maintaining high accuracy. However, including MLPs raised usage to 60.71%, indicating that while attention sparsity is possible, full circuit function often relies on broader model interaction.
@@ 6 Two-Phase SQL Generation: Intent First, Grounding Later
Figure 3: Emergence of SQL intent (via keywords) precedes table name resolution during generation in BM2-CS3. LogitLens reveals elevated logit probabilities for SQL keywords early, followed by table tokens at deeper layers.
@@ 8 Conclusion
Acknowledgements
We thank Narmeen Oozeer for her invaluable early feedback on the project and for proposing Edge Attribution Patching as a linear approximation of the ACDC idea. We are also grateful to Shriyash Upadhyay for his thoughtful feedback and assistance with communication around the project. We appreciate Antia Garcia Casal for her help with figures and for shaping the overall style of our visualizations.
@@ G.1.2 EAP results
For each run, we select the top 10 edges per batch and retain only the edges that appeared at least T% of the time across batches. For CS1 and CS2, we set T = 100, while for CS3, we set T = 80.